# Chenxu Wang

### v4

@@ 3.3 Ablation experiment for YZS-Model
In an effort to better understand the sources of the interpretability performance of the YZS-Model, we carried out a series of ablation experiments. Ablation experiments operate by sequentially controlling individual conditions or parameters and observing the resulting performance changes, thereby identifying the contributions of various model components to the overall performance. To explore the model’s dependency on the training dataset, we began by randomly shuffling the dataset and then splitting it into 25%, 50%, and 75% portions of the original dataset size. We trained the model using these subsets, and the results are illustrated in Figure 11.
Figure 11: .
As illustrated in Figure 10, the model’s performance is less than satisfactory when trained on only 25% of the original dataset. Nevertheless, a substantial improvement in prediction accuracy is observed when the training set is increased to 50%. Further expansion of the training data to 75% and 100% results in continued performance enhancement, though the improvement plateaus beyond the 75% threshold. This demonstrates that our model effectively utilizes training data, achieving baseline performance with just half of the original dataset.
@@
To gain further insights into the contributions of different modules of the YZS-Model during prediction, we performed a series of combinatorial module ablation experiments. We systematically combined the three crucial sub-modules in various ways and conducted ablation tests on these configurations. This approach allows us to evaluate the impact of each module on the prediction accuracy, thereby deepening our understanding of the model. The results of these experiments are illustrated in Figure 11.
Firstly, we examined the impact of ablating individual modules. By removing the GCN, Transformer, and LSTM modules separately, we observed that eliminating the GCN module had minimal effect on prediction performance, while the removal of the Transformer and LSTM modules resulted in significant performance degradation. Further analysis of dual module ablation experiments revealed that the model achieved optimal performance when both GCN and LSTM were removed, highlighting the substantial positive influence of the Transformer module on prediction results. This finding is corroborated by additional dual module ablation experiments: prediction performance markedly declines when either LSTM or GCN is retained without the Transformer. Thus, we conclude that the Transformer is the critical predictive module in the YZS-Model.
To investigate further the impact of the Transformer’s internal components on prediction results, we performed two sets of ablation experiments. Firstly, we replaced the Transformer’s key components, including the Multi-Head Self-Attention (MSA) mechanism and the Feed-Forward Network (FFN), with identity mappings. Secondly, we conducted a series of ablation experiments to examine how varying the number of layers within the Transformer’s internal components affects prediction performance. The experimental results are shown in Figure 12.
In the first set of experiments, replacing MSA or FFN with identity mappings resulted in a significant increase in RMSE, underscoring the importance of these components to the Transformer. In the second set of layer ablation experiments, RMSE exhibited a trend of initially decreasing and then increasing. We hypothesize that with fewer layers, the Transformer struggles to capture the decisive factors in the data effectively; whereas, with more layers, the complexity of the structure may lead to overfitting, thus increasing RMSE. Notably, although the model performed best with 12 layers on the set2 dataset, its performance was inconsistent across other test sets. We attribute this to the inherent randomness in the training process. Taking all factors into account, we conclude that a 6-layer Transformer is the most appropriate structure for this task.
@@
These experimental findings provide valuable insights into the contributions of each component of the YZS-Model in prediction tasks and offer a foundation for further optimization of the model. Through these ablation experiments, we not only confirmed the critical role of the Transformer module but also revealed the specific impacts of its internal components and layer configurations on model performance, thereby guiding future research and applications.
@@ 4 Conclusion
This research introduces and validates the YZS-Model, a novel deep learning framework that incorporates Graph Convolutional Networks (GCNs), Transformer architectures, and Long Short-Term Memory networks (LSTMs) for accurate solubility prediction of organic drugs. This integration allows the YZS-Model to comprehensively analyze the molecular spatial structures and sequence dynamics, significantly enhancing its predictive performance.
The model has shown superior performance on diverse test sets, particularly excelling in the solubility predictions of anticancer compounds, outperforming existing benchmarks. Moreover, our detailed analyses of feature importance have unveiled critical factors affecting solubility, offering insights crucial for drug design and optimization.
Despite its successes, the model’s performance on molecules with unique and rare structures suggests areas for improvement. Future studies will aim to extend the model’s framework to include techniques such as semi-supervised and unsupervised learning, utilizing unlabeled data to better generalize to unknown datasets.
In summary, the YZS-Model not only showcases its strong potential in solubility prediction but also opens new avenues for deploying deep learning in drug discovery and development. Such technologies could potentially assist in preclinical drug screening and optimizing candidate drugs to enhance bioavailability. Furthermore, they hold promise for integrating personalized genetic information to predict drug responses, thereby informing personalized treatment strategies. We anticipate future endeavors to further explore and extend these innovative technologies to contribute significantly to human health.
Despite our achievements, it is vital to acknowledge that drug development involves a wide array of complex factors beyond the explanatory power of a single model. Therefore, ongoing exploration and enhancement by future researchers are essential to fully exploit the extensive capabilities of deep learning in drug development.

### v5

@@ 3.3 Ablation experiment for YZS-Model
3.3.1 Evaluation of Training Data Dependency
To further clarify the sources of the model’s interpretability performance, we performed a series of ablation experiments on the YZS-Model. The core principle of ablation experiments is to systematically control individual conditions or parameters and observe the changes in outcomes, thus revealing the contribution of different parts of the model to its overall performance. Firstly, to assess the model’s dependence on the training set, we randomly shuffled the training dataset and then sequentially divided it into 25%, 50%, and 75% of its original volume. We trained the model using these subsets, and the results are illustrated in Figure 12.
Figure 12: Performance results of the YZS-Model under different training set sizes in ablation studies.
Figure 12. illustrates the impact of training data size on the model’s RMSE across different datasets. In the ablation experiments, the training data was divided into subsets of 25%, 50%, 75%, and 100% of the original data size. The results indicate that using only 25% of the training data leads to poor performance. However, increasing the training data size to 50% significantly improves the predictive accuracy. Further increasing the data size to 75% and 100% continues to enhance the prediction results, although the performance gains beyond 75% are marginal. This demonstrates that the model is highly dependent on the amount of training data, achieving baseline performance with just 50% of the original training data.
As depicted in Figure 10, the model’s performance is unsatisfactory when only 25% of the original dataset is used. However, when the training set size is increased to 50%, there is a notable improvement in prediction accuracy. Further increasing the training set to 75% and 100% continues to enhance the prediction results, although the performance gains become less significant after 75%. This suggests that our model has a robust dependence on training data, achieving baseline performance with just 50% of the original training data.
3.3.2 Contribution Evaluation of YZS-Model Modules
@@
To further explore the contributions of different modules in the YZS-Model during prediction, we conducted combination module ablation experiments. Specifically, we combined three essential submodules in various ways and performed ablation experiments on these combinations. This method allows us to determine the contribution of each module to the prediction results, leading to a deeper understanding of the model. The experimental results are presented in Figure 13.
Figure 13: Performance results of the YZS-Model under different module combinations in ablation studies. The figure compares the RMSE of different models across various datasets. In the core component removal experiments shown on the left, the GCN, Transformer, and LSTM modules were individually removed and compared with the original model. The results indicate that removing the Transformer and LSTM modules has a significant impact on model performance, while the removal of the GCN module has a smaller effect. On the right, the single module experiments demonstrate that models using only LSTM, Transformer, or GCN individually perform worse than the original model, further confirming the importance of the synergistic effect of these modules in enhancing model performance.
The figure compares the RMSE of different models across various datasets. In the core component removal experiments shown on the left, the GCN, Transformer, and LSTM modules were individually removed and compared with the original model. The results indicate that removing the Transformer and LSTM modules has a significant impact on model performance, while the removal of the GCN module has a smaller effect. On the right, the single module experiments demonstrate that models using only LSTM, Transformer, or GCN individually perform worse than the original model, further confirming the importance of the synergistic effect of these modules in enhancing model performance.
Initially, by examining the performance of the original model, we observe that the complete model achieves an RMSE of 0.57, significantly outperforming other configurations, indicating that the model’s integrity is critical for prediction performance. Next, we analyzed the ablation effects of individual modules. By separately removing the GCN, Transformer, and LSTM modules, we found that the exclusion of the GCN module had a minor impact on prediction performance, whereas the removal of the Transformer and LSTM modules had a more substantial effect. This suggests that while the GCN is involved in feature extraction, it has a lesser impact on prediction performance compared to the more crucial roles of the Transformer in capturing complex feature relationships and the LSTM in uncovering deep sequential dependencies within molecular sequences. Further analysis of dual-module ablation experiments indicates that the model performs best when both the GCN and LSTM are excluded, highlighting the significant positive influence of the Transformer module on prediction results. This conclusion is corroborated by other dual-module ablation experiments: prediction performance significantly deteriorates when the Transformer is absent but the LSTM or GCN are present. Therefore, we can conclude that the Transformer is the key predictive module in the YZS-Model.
3.3.3 Detailed Evaluation of Transformer Module
To further investigate the contributions of the internal components of the Transformer model to predictive accuracy, and to elucidate its internal mechanisms and the importance of its components, we conducted two types of ablation experiments. The primary objective was to identify and quantify the impact of key components—Multi-Head Self-Attention (MSA) and Feed-Forward Neural Network (FFN)—as well as the number of layers on overall model performance. We employed the following methods.
Core Component Ablation: In the first set of experiments, we replaced core components of the Transformer—MSA and FFN—with identity mappings. Specifically, the MSA mechanism was substituted with an identity mapping, allowing inputs to bypass the self-attention computation and directly pass to the next layer. Similarly, the FFN was replaced with an identity mapping, permitting inputs to proceed to the subsequent layer without undergoing the feed-forward transformation. These modifications aimed to assess the importance of MSA and FFN in learning complex features and enhancing predictive accuracy.
Layer Ablation: In the second set of experiments, we varied the number of Transformer layers to observe the effects of different model depths on prediction performance. We tested configurations with reduced layers, specifically architectures with 2, 4, and 6 layers, and configurations with increased layers, specifically architectures with 8, 10, and 12 layers. The goal was to understand how the number of layers influences the model’s ability to capture decisive factors in the data and balance between model complexity and overfitting.
Figure 14: Performance results of the YZS-Model under ablation of specific internal modules of the Transformer. Figure illustrates the impact of replacing core components and adjusting the number of layers on the model’s RMSE across different datasets. In the core component ablation experiments, the Multi-Head Self-Attention (MSA) and Feed-Forward Neural Network (FFN) were replaced with identity mappings. The results indicate that both MSA and FFN are crucial to the Transformer’s attention mechanism. In the layer ablation experiments, Transformer architectures with 2, 4, 6, 8, 10, and 12 layers were tested. It was found that the 6-layer Transformer consistently performed the best across multiple datasets.
As figure 14, in the core component ablation experiments, replacing MSA or FFN with identity mappings resulted in a significant increase in RMSE, underscoring the crucial role of these components in the Transformer’s attention mechanism. Specifically, replacing MSA led to a marked increase in RMSE, highlighting its essential function in capturing intricate relationships within the input data. Similarly, replacing FFN resulted in a significant RMSE increase, emphasizing its vital role in processing and transforming feature representations. In the layer ablation experiments, RMSE exhibited a trend of initially decreasing and then increasing with the number of layers. Models with 2 and 4 layers had high RMSE, suggesting underfitting due to inadequate capacity to capture key features. The 6-layer Transformer achieved the lowest RMSE, indicating an optimal balance where the model sufficiently captured features without overfitting. However, models with 8, 10, and 12 layers showed increased RMSE, suggesting overfitting due to excessive complexity. Notably, the 12-layer Transformer performed best on the set2 dataset but exhibited instability across other test datasets, likely due to randomness in the training process. Overall, the results identified the 6-layer Transformer as the most suitable architecture for this task, offering an optimal trade-off between model complexity and generalization ability.
@@ 3.3.3 Detailed Evaluation of Transformer Module
These experimental results provide valuable insights into the contributions of various components within the YZS-Model to prediction tasks, forming a basis for further model optimization. The ablation experiments not only confirmed the critical role of the Transformer module but also highlighted the specific effects of its internal components and layer number on model performance, guiding future research and applications.
@@ 4 Conclusion
In this study, we developed and validated the YZS-Model, an innovative deep learning framework based on Graph Convolutional Networks (GCNs), Transformer architecture, and Long Short-Term Memory networks (LSTMs), designed for the precise prediction of organic drug solubility. By integrating these three advanced deep learning technologies, the YZS-Model can deeply analyze the spatial structure and sequence dynamics of drug molecules, resulting in significant performance improvements in the prediction of organic drug solubility.
Our model has demonstrated superior performance across different test sets, particularly excelling in solubility predictions for anticancer compounds, where it outperformed existing benchmark models in terms of accuracy and generalization capability. Additionally, through a series of detailed feature importance analyses, we have identified key factors affecting drug molecule solubility, providing valuable insights for drug design and optimization.
However, we also observed that the model’s accuracy could be further improved when dealing with molecules with specific and rare structural features. To address this challenge, future research will consider further expanding the model by exploring other machine learning strategies, including semi-supervised and unsupervised learning methods, to utilize unlabeled data for model training, thereby enhancing the model’s generalization ability to unknown datasets. We also plan to introduce multi-scale attention mechanisms to more precisely capture the complex interactions within molecules, improving prediction accuracy. Additionally, ensemble learning strategies will be employed to further enhance prediction performance.
In summary, the YZS-Model not only demonstrates strong potential in predicting the solubility of organic drugs but also paves new avenues for the application of deep learning technologies in drug discovery and development. The application of deep learning technologies is expected to assist in preclinical drug screening and optimize the chemical properties of candidate drugs to increase their bioavailability. Moreover, in the field of personalized medicine, this model holds promise for integrating individual genetic backgrounds to predict drug responses, thereby informing the development of personalized treatment strategies. We anticipate that future research will further explore and expand these technologies, making greater contributions to human health.
Despite our achievements, it is essential to recognize that drug development is a complex and dynamic process, involving factors far beyond the explanatory power of a single model. Therefore, we encourage ongoing exploration and improvement by future researchers to fully leverage the vast potential of deep learning in drug development.